# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ A.7.3 KSD goodness-of-testing testing
$\big{\|}{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ c\big{\|}{\bm{\psi}}\big{\|}_{L^{2}}^{2}+\widetilde{C}\sum_{i=1}^{d}\lambda^{2s}_{i}.$ (155)
Then, the overall uniform separation rate of Equation 153 becomes
$\big{\|}{\bm{\psi}}\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sum_{i=1}^{d}\lambda^{2s}_{i}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (156)
@@
$\big{\|}{\bm{\psi}}\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N}}\right)^{2s/(4s+5d)}$ (157)
@@ A.8 Proof sketch of efficient L2 separation
We detail the proof structure of the efficient $L^{2}$ separation results: Equations 23 and 38 proved in Schrab et al. (2022b, Theorem 1), and Equation 51 proved here.
@@
As in the case of kernel separation (Section A.1), the proof of $L^{2}$ separation in Section A.7 relies on two exponential concentration results: one for the test statistic (Equation 105), and one for the bootstrapped statistic leading to a quantile bound (Equation 111). To prove the desired efficient $L^{2}$ separation rates of Equations 23, 38 and 51, it then suffices to derive versions of the results of Equations 105 and 111 with $N$ replaced by $|{\mathcal{D}}|/N$. We now illustrate how this can be done.
An equivalent version to the exponential concentration result of Equation 105 for incomplete U-statistics is provided by Maurer (2022, Theorem 3.3) guaranteeing that, with probability at least $21-\beta/$ 2, we have [^11] Referring to the notation of Maurer (2022, Theorem 3.3), we have $\alpha,\beta,\gamma$ bounded by constants, $A\leq N/|{\mathcal{D}}|$, $B\leq N^{2}/|{\mathcal{D}}|^{2}$, $C\leq N/|{\mathcal{D}}|$, where the bound for $A$ holds assuming that the number of entries of $\mathcal{D}$ in each row of the $N\times N$ kernel/core matrix is at most of the order of $\sqrt{|{\mathcal{D}}|}$. Intuitively this requires that the entries of ${\mathcal{D}}$ are spread out around the kernel matrix, they cannot all be concentrated on a same row when $|{\mathcal{D}}|$ is small compared to $N^{2}$.
$\big{|}U_{0}-{\mathrm{Kdisc}}^{2}\big{|}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \sqrt{\frac{N}{|{\mathcal{D}}|}\sigma_{1}^{2}\log\biggl{(}\frac{1}{\beta}\biggr{)}}+{\frac{N}{|{\mathcal{D}}|}\log\biggl{(}\frac{1}{\beta}\biggr{)}}$ (158)
where $\sigma_{1}^{2}\coloneqq{\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}$ as in Equation 106.
To derive an equivalent version of Equation 111, we note that the exponential concentration bound for i.i.d. Rademacher chaos of de la Peña and Giné (1999b, Corollary 3.2.6) can be applied to the case of incomplete U-statistics (see Schrab et al., 2022b, Theorem 1) to obtain that, with probability at least $21-\beta/$ 2, we have
$q_{1-\alpha}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \sqrt{\frac{1}{|{\mathcal{D}}|^{2}}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}}\log\!\left(\frac{1}{\alpha}\right)\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{N}{|{\mathcal{D}}|}\frac{1}{\sqrt{\lambda_{1}\cdots\lambda_{d}}}\sqrt{\log\!\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (159)
using the bound of Equation 108 relying on Bernstein’s inequality, where $q_{1-\alpha}$ is the $(1-\alpha)$-quantile of the incomplete bootstrapped U-statistics.
With these two exponential concentration bounds adapted for incomplete U-statistics, the efficient $L^{2}$ separation results of Equations 23, 38 and 51 can be proved by following the exact same resoning presented in Section A.7.
@@ A.9 Proof sketch of aggregated L2 separation
We detail the proof structure of the aggregated $L^{2}$ separation results: Equation 22 proved in Schrab et al. (2023, Corollary 10), Equation 37 proved in Schrab et al. (2022b, Theorem 3) with estimated quantiles (and in Albert et al., 2022, Corollary 3 with theoretical quantiles), and Equation 50 proved here.
@@
Let $\lambda_{1}=\dots=\lambda_{d}=\lambda$, and consider the aggregated test (Section 2) over the bandwidths
$\Lambda\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ \left\{2^{-\ell}:\ell\in\left\{1,\dots,\left\lceil\frac{2}{d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil\right\}\right\},$ (160)
@@
~$\widetilde{\ell}\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ \left\lceil\frac{2}{4s+d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left\lceil\frac{2}{d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil$ (161)
@@
$\frac{1}{2}\!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \widetilde{\lambda}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}.$ (162)
@@
$\widetilde{\lambda}\leavevmode\nobreak\ \asymp\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}$ (163)
like the optimal bandwidth in Section A.7. Then, as in Equation 117, the uniform separation rate of this specific test, with bandwith $\widetilde{\lambda}=2^{-\widetilde{\ell}}$ and adjusted level $\alpha w_{\widetilde{\ell}}$, is
$\big{\|}\psi\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sum_{i=1}^{d}\widetilde{\lambda}^{2s}+\frac{1}{N{\widetilde{\lambda}^{d/2}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha w_{\widetilde{\ell}}}\right)$ (164)
@@
$\big{\|}\psi\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sum_{i=1}^{d}\widetilde{\lambda}^{2s}+\frac{\log(\log(N))}{N{\widetilde{\lambda}^{d/2}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (165)
since $w_{\widetilde{\ell}}={6}/{{\widetilde{\ell}}^{2}\pi^{2}}$ giving $\ln\!\left({1}/{w_{\widetilde{\ell}}}\,\right)\lesssim\ln(\widetilde{\ell}\,)\lesssim\ln(\ln(N)).$ Substituting the expression of Equation 163 for the specific bandwidth $\widetilde{\lambda}$ of the collection $\Lambda$, we get that the overall aggregated test over $\Lambda$ controls the type II error by $\beta$ whenever
$\big{\|}\psi\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}$ (166)
for any Sobolev smoothness $s>0$ for MMD and KSD, and for any Sobolev smoothness $s\geq(d_{\mathcal{X}}+d_{\mathcal{Y}})/4$ for HSIC (as in Section A.7).
@@ A.10 Proof sketch of aggregated efficient L2 separation
By combining the reasoning of the aggregated and efficient $L^{2}$ separation results in Sections A.9 and A.8, respectively, one obtains the aggregated efficient $L^{2}$ separation results: Equation 24 proved in Schrab et al. (2022b, Theorem 2), Equation 39 proved in Schrab et al. (2022b, Theorem 2), and Equation 52 proved here.
@@ A.11 Proof sketch of differentially private L2 separation
We detail the proof structure of the differentially private $L^{2}$ separation results: Equations 25, 26 and 27 proved in Kim and Schrab (2023, Theorem 9), and Equations 40, 41 and 42 proved in Kim and Schrab (2023, Theorem 14).
The proof structure is similar to the one of the non-private case depicted in Section A.7 with two major differences. The first one is that the results need to be adapted from holding for U-statistics to holding for V-statistics, this is done by expressing the V-statistics in terms of the U-statistics (Kim and Schrab, 2023, Equation 38 and Lemma 22). The second difference is that the added privatisation Laplacian noise needs to be taken into account, it is scaled by the global sensivity of the square-rooted V-statistic which is of order $1/(N\!\sqrt{\lambda_{1}\cdots\lambda_{d}})$. For detailed proofs, see Kim and Schrab (2023, Theorems 9 and 14), in these the dependence on $\beta$ is polynomial, we believe it can be improved to be logarithmic by relying on Berstein’s inequality as done in Section A.7.

### v2

@@ A.7.3 KSD goodness-of-testing testing
$\big\|{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big\|_{L^{2}}^{2}\penalty 10000\ \leq\penalty 10000\ c\big\|{\bm{\psi}}\big\|_{L^{2}}^{2}+\widetilde{C}\sum_{i=1}^{d}\lambda^{2s}_{i}.$ (155)
Then, the overall uniform separation rate of Equation˜153 becomes
$\big\|{\bm{\psi}}\big\|_{L^{2}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \sum_{i=1}^{d}\lambda^{2s}_{i}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (156)
@@
$\big\|{\bm{\psi}}\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N}}\right)^{2s/(4s+5d)}$ (157)
@@ A.8 Proof sketch of efficient L2 separation
We detail the proof structure of the efficient $L^{2}$ separation results: Equations˜23 and 38 proved in Schrab et al. (2022b, Theorem 1), and Equation˜51 proved here.
@@
As in the case of kernel separation (Section˜A.1), the proof of $L^{2}$ separation in Section˜A.7 relies on two exponential concentration results: one for the test statistic (Equation˜105), and one for the bootstrapped statistic leading to a quantile bound (Equation˜111). To prove the desired efficient $L^{2}$ separation rates of Equations˜23, 38 and 51, it then suffices to derive versions of the results of Equations˜105 and 111 with $N$ replaced by $|{\mathcal{D}}|/N$. We now illustrate how this can be done.
An equivalent version to the exponential concentration result of Equation˜105 for incomplete U-statistics is provided by Maurer (2022, Theorem 3.3) guaranteeing that, with probability at least $1-\beta/2$, we have [^11] Referring to the notation of Maurer (2022, Theorem 3.3), we have $\alpha,\beta,\gamma$ bounded by constants, $A\leq N/|{\mathcal{D}}|$, $B\leq N^{2}/|{\mathcal{D}}|^{2}$, $C\leq N/|{\mathcal{D}}|$, where the bound for $A$ holds assuming that the number of entries of $\mathcal{D}$ in each row of the $N\times N$ kernel/core matrix is at most of the order of $\sqrt{|{\mathcal{D}}|}$. Intuitively this requires that the entries of $\mathcal{D}$} are spread out around the kernel matrix, they cannot all be concentrated on a same row when $|{\mathcal{D}}|$ is small compared to $N^{2}$.
$\big|U_{0}-{\mathrm{Kdisc}}^{2}\big|\penalty 10000\ \lesssim\penalty 10000\ \sqrt{\frac{N}{|{\mathcal{D}}|}\sigma_{1}^{2}\log\biggl(\frac{1}{\beta}\biggr)}+{\frac{N}{|{\mathcal{D}}|}\log\biggl(\frac{1}{\beta}\biggr)}$ (158)
where $\sigma_{1}^{2}\coloneqq{\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}$ as in Equation˜106.
To derive an equivalent version of Equation˜111, we note that the exponential concentration bound for i.i.d. Rademacher chaos of de la Peña and Giné (1999b, Corollary 3.2.6) can be applied to the case of incomplete U-statistics (see Schrab et al., 2022b, Theorem 1) to obtain that, with probability at least $1-\beta/2$, we have
$q_{1-\alpha}\penalty 10000\ \lesssim\penalty 10000\ \sqrt{\frac{1}{|{\mathcal{D}}|^{2}}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}}\log\!\left(\frac{1}{\alpha}\right)\penalty 10000\ \lesssim\penalty 10000\ \frac{N}{|{\mathcal{D}}|}\frac{1}{\sqrt{\lambda_{1}\cdots\lambda_{d}}}\sqrt{\log\!\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (159)
using the bound of Equation˜108 relying on Bernstein’s inequality, where $q_{1-\alpha}$ is the $(1-\alpha)$-quantile of the incomplete bootstrapped U-statistics.
With these two exponential concentration bounds adapted for incomplete U-statistics, the efficient $L^{2}$ separation results of Equations˜23, 38 and 51 can be proved by following the exact same resoning presented in Section˜A.7.
@@ A.9 Proof sketch of aggregated L2 separation
We detail the proof structure of the aggregated $L^{2}$ separation results: Equation˜22 proved in Schrab et al. (2023, Corollary 10), Equation˜37 proved in Schrab et al. (2022b, Theorem 3) with estimated quantiles (and in Albert et al., 2022, Corollary 3 with theoretical quantiles), and Equation˜50 proved here.
@@
Let $\lambda_{1}=\dots=\lambda_{d}=\lambda$, and consider the aggregated test (Section˜2) over the bandwidths
$\Lambda\penalty 10000\ \coloneqq\penalty 10000\ \left\{2^{-\ell}:\ell\in\left\{1,\dots,\left\lceil\frac{2}{d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil\right\}\right\},$ (160)
@@
$\widetilde{\ell}\penalty 10000\ \coloneqq\penalty 10000\ \left\lceil\frac{2}{4s+d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil\penalty 10000\ \leq\penalty 10000\ \left\lceil\frac{2}{d}\log_{2}\!\left(\frac{N/\log(\log(N))}{\log(1/\alpha)\log(1/\beta)}\right)\right\rceil$ (161)
@@
$\frac{1}{2}\!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}\penalty 10000\ \leq\penalty 10000\ \widetilde{\lambda}\penalty 10000\ \leq\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}.$ (162)
@@
$\widetilde{\lambda}\penalty 10000\ \asymp\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2/(4s+d)}$ (163)
like the optimal bandwidth in Section˜A.7. Then, as in Equation˜117, the uniform separation rate of this specific test, with bandwith $\widetilde{\lambda}=2^{-\widetilde{\ell}}$ and adjusted level $\alpha w_{\widetilde{\ell}}$, is
[…]
@@
$\big\|\psi\big\|_{L^{2}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \sum_{i=1}^{d}\widetilde{\lambda}^{2s}+\frac{\log(\log(N))}{N{\widetilde{\lambda}^{d/2}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (165)
since $w_{\widetilde{\ell}}={6}/{{\widetilde{\ell}}^{2}\pi^{2}}$ giving $\ln\!\left({1}/{w_{\widetilde{\ell}}}\,\right)\lesssim\ln(\widetilde{\ell}\,)\lesssim\ln(\ln(N)).$ Substituting the expression of Equation˜163 for the specific bandwidth $\widetilde{\lambda}$ of the collection $\Lambda$, we get that the overall aggregated test over $\Lambda$ controls the type II error by $\beta$ whenever
$\big\|\psi\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}$ (166)
for any Sobolev smoothness $s>0$ for MMD and KSD, and for any Sobolev smoothness $s\geq(d_{\mathcal{X}}+d_{\mathcal{Y}})/4$ for HSIC (as in Section˜A.7).
@@ A.10 Proof sketch of aggregated efficient L2 separation
By combining the reasoning of the aggregated and efficient $L^{2}$ separation results in Sections˜A.9 and A.8, respectively, one obtains the aggregated efficient $L^{2}$ separation results: Equation˜24 proved in Schrab et al. (2022b, Theorem 2), Equation˜39 proved in Schrab et al. (2022b, Theorem 2), and Equation˜52 proved here.
@@ A.11 Proof sketch of differentially private L2 separation
We detail the proof structure of the differentially private $L^{2}$ separation results: Equations˜25, 26 and 27 proved in Kim and Schrab (2023, Theorem 9), and Equations˜40, 41 and 42 proved in Kim and Schrab (2023, Theorem 14).
The proof structure is similar to the one of the non-private case depicted in Section˜A.7 with two major differences. The first one is that the results need to be adapted from holding for U-statistics to holding for V-statistics, this is done by expressing the V-statistics in terms of the U-statistics (Kim and Schrab, 2023, Equation 38 and Lemma 22). The second difference is that the added privatisation Laplacian noise needs to be taken into account, it is scaled by the global sensivity of the square-rooted V-statistic which is of order $1/(N\!\sqrt{\lambda_{1}\cdots\lambda_{d}})$. For detailed proofs, see Kim and Schrab (2023, Theorems 9 and 14), in these the dependence on $\beta$ is polynomial, we believe it can be improved to be logarithmic by relying on Berstein’s inequality as done in Section˜A.7.